# Restricted Maximum Likelihood Estimation in Generalized Linear Mixed Models

### v1

@@ Abstract
Restricted maximum likelihood (REML) estimation is a widely accepted and frequently used method for fitting linear mixed models, with its principal advantage being that it produces less biased estimates of the variance components. However, the concept of REML does not immediately generalize to the setting of non-normally distributed responses, and it is not always clear the extent to which, either asymptotically or in finite samples, such generalizations reduce the bias of variance component estimates compared to standard unrestricted maximum likelihood estimation. In this article, we review various attempts that have been made over the past four decades to extend REML estimation in generalized linear mixed models. We establish four major classes of approaches, namely approximate linearization, integrated likelihood, modified profile likelihoods, and direct bias correction of the score function, and show that while these four classes may have differing motivations and derivations, they often arrive at a similar if not the same REML estimate. We compare the finite sample performance of these four classes through a numerical study involving binary and count data, with results demonstrating that they perform similarly well in reducing the finite sample bias of variance components.
Keywords: bias correction; modified profile likelihood; profile score function; random effects; variance estimation
@@ 1 Introduction
Historically, REML estimation is often attributed to the seminal contributions of Patterson and Thompson, (1971, 1974), although developments had been made for some balanced analysis of variance (ANOVA) models (Anderson and Bancroft,, 1952; Russell and Bradley,, 1958), and then ‘all’ balanced ANOVA models by Thompson, (1962). Following the work of Patterson and Thompson, (1971, 1974), the restricted likelihood function was defined as the marginal likelihood of a set of residual contrasts, from which the idea of REML estimation for LMMs arises. Harville, (1974) formulated an alternative derivation of REML based on integrated likelihood using Bayesian arguments, while Cooper and Thompson, (1977) provided another derivation focusing on the autoregressive-moving average model. Laird and Ware, (1982) showed that in balanced ANOVA models, REML estimation is equivalent to maximizing the likelihood obtained by integrating out the fixed effects using a diffuse prior. Another derivation of REML for LMMs comes from Verbyla, (1990), who presented a conditional derivation related to that of Harville, (1974); see Section 2.1. For LMMs specifically, REML was further shown to be a special case of approximate conditional likelihood (Cox and Reid,, 1987), adjusted profile likelihood (McCullagh and Tibshirani,, 1990) and modified profile likelihood (Chapter 9, Severini,, 2000). Ultimately, all these derivations of REML for LMMs arrive at the same objective function, and hence the same estimators of the random effects covariance matrix.
@@
Despite the extensive literature on REML estimation in LMMs, it is not immediately obvious how to define REML outside this class of models, since zero-mean residual contrasts generally do not exist in nonlinear models (Smyth and Verbyla,, 1996). One consequence, and the main focus of this review, is that there have been many attempts to extend REML estimation to generalized linear mixed models (GLMMs) with non-normally distributed responses and a non-identity link function (see, McCulloch and Searle,, 2004, for an overview of GLMMs). Unlike their linear counterparts where the same restricted likelihood function is ultimately reached, with GLMMs different formulations of REML can potentially lead to different objective functions and thus seemingly different REML estimators exhibiting varying empirical behaviour. REML estimation for GLMMs and non-normally distributed response models remains a subject of active research (e.g., Tawiah et al.,, 2019, McNeish,, 2019, Elff et al.,, 2021) while REML methods continue to be applied in various contexts such as bias adjustment for the asymptotic covariance of LMM and GLMM parameters (e.g., Kenward and Roger,, 1997).
@@
In this article, we aim to provide an overview of the various REML estimation methods that have been proposed for GLMMs. We categorize these methods into four main classes of REML approaches: approximate linearization (the largest category), integrated likelihood, modified profile likelihood and direct bias correction of the score function methods. A central result we find is that while these four classes may have differing motivations and algorithmic implementations, they often arrive at a similar if not the same REML estimate. A small simulation study comparing several currently available methods for REML estimation in R for GLMMs with binary and count data reveals that most REML methods perform similarly well in reducing the finite sample bias of the variance components.
@@
The present literature on GLMMs lacks a comprehensive and up-to-date review of the various approaches for REML estimation. Partial summaries of REML procedures have been presented in the past, but these either focus on a particular response type such as binary GLMMs (Reichert,, 1993; Meza et al.,, 2009), or are tailored to particular classes of mixed models such as hierarchical generalized linear models (Lee and Nelder,, 2003; Noh and Lee,, 2007). We seek to fill this gap and provide a overarching review with a particular emphasis on methods that are applicable to a broad range of GLMMs.
@@
In Section 2 we define GLMMs and one approach to constructing REML estimation for LMMs. We then categorize the different methods for REML estimation in GLMMs into four main approaches (based on approximate linearization, integrated likelihood, modified profile likelihood or direct bias adjustment for the profile score function) and describe each approach in its own section (Sections 36). Section 7 provides a numerical study assessing the performance of some REML estimation methods currently available and/or easily implementable in R. Finally, Section 8 offers some concluding remarks and avenues for future research. As a final introductory remark, we note the first letter of the REML acronym has been used to refer to both “restricted” and “residual” maximum likelihood estimation. Historically, these two terms have been employed interchangeably, although in this paper we will use the term “restricted” as it has found more widespread use in the literature.
@@ 2 Generalized Linear Mixed Models
Generalized linear mixed models, which encompass the special case of linear mixed models with normally distributed responses and an identity link, express the mean of the response distribution as a function of fixed and random effects. A general formulation of GLMMs is given by Hui, (2021) among others, where for the $i$th of $n$ observations we have
$y_{i}|\boldsymbol{u}\sim p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi),\quad g(\mu_{i})=\eta_{i}=\boldsymbol{x}_{i}^{\top}\boldsymbol{\beta}+\boldsymbol{z}_{i}^{\top}\boldsymbol{u},\quad\boldsymbol{u}\sim p(\boldsymbol{u};\boldsymbol{\theta})=N\{\boldsymbol{0},\boldsymbol{D}(\boldsymbol{\theta})\},$ (1)
@@
where $y_{i}$ denotes a univariate response variable, $\boldsymbol{x}_{i}$ and $\boldsymbol{z}_{i}$ denote vectors of $p$ fixed and $q$ random effect covariates, respectively, and $\boldsymbol{\beta}$ and $\boldsymbol{u}$ denote the corresponding fixed and random effect coefficients. Conditional on the random effects, the $y_{i}$’s are assumed to be independent observations from an exponential family of distributions $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ with mean $\mbox{E}(y_{i}|\boldsymbol{u})=\mu_{i}$ and dispersion parameter $\phi>0$, where the latter appears in the conditional variance of the distribution, $\mbox{Var}(y_{i}|\boldsymbol{u})=\phi V(\mu_{i})$. The mean $\mu_{i}$ is related to the fixed and random effect covariates via a known link function $g(\cdot)$.
@@
Linear mixed models arise as a special case of (1) when $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ is a normal distribution, $g(\mu)=\mu$ is set to the identity link, and $\phi$ is the error variance such that $V(\mu)=1$. Two prominent non-normal cases of GLMMs are binomial GLMMs where $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ is the binomial distribution, $g(\mu)=\log\{\mu/(1-\mu)\}$ is the canonical logit link, $\phi=1$ is known and $V(\mu)=\mu(1-\mu)$, and Poisson GLMMs where $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ is the Poisson distribution, $g(\mu)=\log(\mu)$ is the canonical log link, $\phi=1$ is known and $V(\mu)=\mu$. The above formulation can be further expanded to involve more than one dispersion parameter, e.g., for Tweedie responses, and the REML methods reviewed in this article can be modified to accommodate such settings. However, for ease of presentation we will focus on the above formulation and refer the reader to McCullagh and Nelder, (1989) and Wood, (2017) for general overviews on exponential families of distributions for regression modeling.
@@
We assume the random effects in (1) are drawn from a multivariate normal distribution with a mean vector of zeros, and a random effects covariance matrix $\mbox{Cov}(\boldsymbol{u})=\boldsymbol{D}(\boldsymbol{\theta})$ which depends on a set of parameters $\boldsymbol{\theta}$. While it is possible to consider non-normally distributed random effects (see Section S.1 of the supplementary material for a discussion on REML estimation in hierarchical generalized linear models), we restrict this review to the normal random effects case given this is the most prevalent assumption in practice (Breslow and Clayton,, 1993; McCulloch and Searle,, 2004; Stroup,, 2012). Note $\boldsymbol{D}(\boldsymbol{\theta})$ can and often does possess a large amount of structure, e.g., it may have a block-diagonal or fully diagonal structure in the case of independent cluster GLMMs. The diagonal elements of $\boldsymbol{D}(\boldsymbol{\theta})$ are often referred to as variance components, although in this article we will use the terms “variance components” and “random effects covariance matrix” interchangeably to mean $\boldsymbol{\theta}$ or $\boldsymbol{D}(\boldsymbol{\theta})$.
@@
Next, we present a vectorized version of the GLMM in (1). Define $\boldsymbol{y}=(y_{1},\ldots,y_{n})^{\top}$, $\boldsymbol{\mu}=(\mu_{1},\ldots,\mu_{n})^{\top}$, $\boldsymbol{\eta}=(\eta_{1},\ldots,\eta_{n})^{\top}$, and let $\boldsymbol{X}$ and $\boldsymbol{Z}$ denote $n\times p$ and $n\times q$ matrices formed by stacking the $\boldsymbol{x}_{i}$’s and $\boldsymbol{z}_{i}$’s as row vectors, respectively. We can then write the mean model of the GLMM as $g(\boldsymbol{\mu})=\boldsymbol{\eta}=\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u}$, where the link function is applied element-wise to $\boldsymbol{\mu}$. This vectorized form of the mean model shows that equation (1) is broad enough to accommodate many applications of GLMMs. For instance, for independent clustered, nested or multi-level GLMMs, the columns of $\boldsymbol{Z}$ can be rearranged so that $\boldsymbol{Z}=[\boldsymbol{Z}_{1}\ldots\boldsymbol{Z}_{K}]$, with all the submatrices $\boldsymbol{Z}_{k}$, $k=1,\ldots,K$, being block-diagonal in form, while $\boldsymbol{D}(\boldsymbol{\theta})$ is also written in a block-diagonal form. Crossed-effect GLMMs are a case where the submatrices $\boldsymbol{Z}_{k}$ are not simultaneously block-diagonal. In this article, we will as much as possible utilize this broad formulation of GLMMs to keep the review general, with specific cases used as necessary, e.g., when particular papers develop REML estimation for a specific case of (1), and also in our simulation study.
@@
Standard unrestricted maximum likelihood estimation is based on optimizing equation (2) or some variation of this such as approximated marginal log-likelihoods obtained via Laplace, quadrature, variational, and Monte-Carlo approximations; see for example, Breslow and Clayton, (1993), Shun and McCullagh, (1995), McCulloch, (1997), Ormerod and Wand, (2012), Hall et al., (2020), and Hui, (2021) and references therein. Often central to fitting GLMMs is the joint log-likelihood function of the responses and random effects
$\ell_{J}(\boldsymbol{\beta},\phi,\boldsymbol{\theta};\boldsymbol{y},\boldsymbol{u})=\log\{p(\boldsymbol{y}|\boldsymbol{u};\boldsymbol{\beta},\phi)\}+\log\{p(\boldsymbol{u};\boldsymbol{\theta})\};$ (3)
in particular the Laplace approximation is implemented around the maximizer of the joint likelihood function (Breslow and Clayton,, 1993; Shun and McCullagh,, 1995).

### v2

@@ Abstract
Restricted maximum likelihood (REML) estimation is a widely accepted and frequently used method for fitting linear mixed models, with its principal advantage being that it produces less biased estimates of the variance components. However, the concept of REML does not immediately generalize to the setting of non-normally distributed responses, and it is not always clear the extent to which, either asymptotically or in finite samples, such generalizations reduce the bias of variance component estimates compared to standard unrestricted maximum likelihood estimation. In this article, we review various attempts that have been made over the past four decades to extend REML estimation in generalized linear mixed models. We establish four major classes of approaches, namely approximate linearization, integrated likelihood, modified profile likelihoods, and direct bias correction of the score function, and show that while these four classes may have differing motivations and derivations, they often arrive at a similar if not the same REML estimate. We compare the finite sample performance of these four classes, along with methods for REML estimation in hierarchical generalized linear models, through a numerical study involving binary and count data, with results demonstrating that all approaches perform similarly well reducing the finite sample size bias of variance components. Overall, we believe REML estimation should more widely adopted by practitioners using generalized linear mixed models, and that the exact choice of which REML approach to use should, at this point in time, be driven by software availability and ease of implementation.
bias correction,
modified profile likelihood,
profile score function,
random effects,
variance estimation,
keywords:
[…]
and
@@ 1 Introduction
Historically, REML estimation is often attributed to the seminal contributions of Patterson and Thompson (1971, 1974), although prior to these, developments had been made for some balanced analysis of variance (ANOVA) models (Anderson and Bancroft, 1952; Russell and Bradley, 1958), and then ‘all’ balanced ANOVA models by Thompson (1962). Following the work of Patterson and Thompson (1971, 1974), the restricted likelihood function was defined as the marginal likelihood of a set of residual contrasts, from which the idea of REML estimation for LMMs arises. Harville (1974) formulated an alternative derivation of REML based on integrated likelihood using Bayesian arguments, while Cooper and Thompson (1977) provided another derivation focusing on the autoregressive-moving average model. Laird and Ware (1982) showed that in balanced ANOVA models, REML estimation is equivalent to maximizing the likelihood obtained by integrating out the fixed effects using a diffuse prior. Another derivation of REML for LMMs comes from Verbyla (1990), who presented a conditional derivation related to that of Harville (1974); see Section 2.1. For LMMs specifically, REML was further shown to be a special case of approximate conditional likelihood (Cox and Reid, 1987), adjusted profile likelihood (McCullagh and Tibshirani, 1990) and modified profile likelihood (Chapter 9, Severini, 2000). Ultimately, all these derivations of REML for LMMs arrive at the same objective function, and hence the same estimators of the random effects covariance matrix.
@@
Despite the extensive literature on REML estimation in LMMs, it is not immediately obvious how to define REML outside this class of models, since zero-mean residual contrasts generally do not exist in nonlinear models (Smyth and Verbyla, 1996). One consequence, and the main focus of this review, is that there have been many attempts to extend REML estimation to generalized linear mixed models (GLMMs) with non-normally distributed responses and a non-identity link function (see, McCulloch and Searle, 2004, for an overview of GLMMs). Unlike their linear counterparts where the same restricted likelihood function is ultimately reached, with GLMMs different formulations of REML can potentially lead to different objective functions and thus seemingly different REML estimators exhibiting varying empirical behaviour.
REML estimation for GLMMs and non-normally distributed response models remains a subject of active research (see, e.g., Tawiah et al., 2019, McNeish, 2019, Elff et al., 2021, and Han and Lee, 2024), while REML methods continue to be applied in various contexts such as model selection (Verbyla, 2019), bias adjustment for the asymptotic covariance of LMM and GLMM parameters (e.g., Kenward and Roger, 1997, and Watson, 2023), meta-analysis (e.g., Bakbergenuly and Kulinskaya, 2018, and Lin and Chu, 2020), or factor analysis (e.g., Noh et al., 2019). More generally, REML has found widespread adoption among practitioners and applied scientists in data analyses involving GLMMs thanks to the availability of software such as the package glmmTMB (Brooks et al., 2017) in R (R Core Team, 2024) and the GLIMMIX procedure in SAS (SAS Institute Inc., 2015). In genome-wide association studies, the usage of REML, e.g., for estimation of genetic correlations, has also become popular thanks to procedures such as BOLT-REML (Loh et al., 2015a) available in the software package BOLT-LMM (Loh et al., 2015b).
@@
In this article, we aim to provide an overview of the various REML estimation methods that have been proposed for GLMMs. We categorize these methods into four main classes of REML approaches: approximate linearization (the largest category), integrated likelihood, modified profile likelihood and direct bias correction of the score function methods. We also review REML estimation for the class of hierarchical generalized linear models (HGLMs), of which GLMMs can be considered a particular class. A central result we find is that while these different approaches to REML estimation may have differing motivations and algorithmic implementations, they often arrive at a similar if not the same estimate. A simulation study comparing several currently available methods for REML estimation in R for GLMMs with binary and count data reveals that most REML methods perform similarly well in reducing the finite sample bias of the variance components, offering noticeable bias reductions, especially when the number of fixed effects is large relative to the number of random effects.
@@
The present literature on GLMMs lacks a comprehensive and up-to-date review of the various approaches for REML estimation. Partial summaries of REML procedures have been presented in the past, but these either focus on a particular response type such as binary GLMMs (Reichert, 1993; Meza, Jaffrézic and Foulley, 2009), or are tailored to particular cases such as HGLMs (Lee and Nelder, 2003; Noh and Lee, 2007). We seek to fill this gap and provide a overarching review with a particular emphasis on methods that are applicable to a broad range of mixed models with non-Gaussian distributed responses.
Our overall conclusions and take-home messages from this review can be summarized as follows. First, we believe the extension of REML to GLMMs should be more widely adopted by practitioners, given our numerical results clearly show that it is worthwhile when the goal is to reduce bias in variance component estimation in GLMMs. Second, given the current theoretical research on REML for GLMMs is rather piecemeal and more motivational than rigorous, we encourage more precise asymptotic investigations into the various flavors of REML estimation for GLMMs, for instance, deriving under what conditions it can remove the leading bias term in the estimation of variance components. Finally, given the above two points, we argue that at this current point in time the choice of which REML implementation to use in mixed model analysis should be driven by practical constraints such as software availability and ease of implementation more than theory.
@@
In Section 2 we define GLMMs and one approach to constructing REML estimation for LMMs. We then categorize the different methods for REML estimation in GLMMs into four main approaches (based on approximate linearization, integrated likelihood, modified profile likelihood or direct bias adjustment for the profile score function) and describe each approach in its own section (Sections 36). Methods developed specifically for the class of HGLMs are reviewed in Section 7. Section 8 provides a numerical study assessing the performance of some REML estimation methods currently available and/or easily implementable in R. Finally, Section 9 summarizes our take-home messages and offers some avenues for future research. As a final introductory remark, we note the first letter of the REML acronym has been used to refer to both “restricted” and “residual” maximum likelihood estimation. Historically, these two terms have been employed interchangeably, although in this paper we will use the term “restricted” as it has found more widespread use in the literature.
@@ 2 Generalized Linear Mixed Models
Generalized linear mixed models, which encompass the special case of linear mixed models with normally distributed responses and an identity link, express the mean of the response distribution as a function of fixed and random effects. A general formulation of GLMMs is given by Hui (2021) among others. For a vector $\boldsymbol{y}=(y_{1},\ldots,y_{n})^{\top}$ of $n$ responses, the corresponding conditional mean vector $\boldsymbol{\mu}=(\mu_{1},\ldots,\mu_{n})^{\top}$ is modeled as
$\begin{array}[]{c}\boldsymbol{y}|\boldsymbol{u}\sim\prod_{i=1}^{n}p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)\\[6.45831pt] g(\boldsymbol{\mu})=\boldsymbol{\eta}=\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u},\\[6.45831pt] \boldsymbol{u}\sim p(\boldsymbol{u};\boldsymbol{\theta})=N\{\boldsymbol{0},\boldsymbol{D}(\boldsymbol{\theta})\},\end{array}$ (1)
@@
where $g(\cdot)$ is a known link function applied element-wise to $\boldsymbol{\mu}$, $\boldsymbol{\eta}=(\eta_{1},\ldots,\eta_{n})^{\top}$ denotes the linear predictor vector, $\boldsymbol{X}$ and $\boldsymbol{Z}$ denote $n\times p$ and $n\times q$ matrices comprising $p$ fixed and $q$ random effect covariates, respectively, and $\boldsymbol{\beta}$ and $\boldsymbol{u}$ denote the corresponding fixed and random effect coefficients. Conditional on the random effects, the $y_{i}$’s are assumed to be independent observations from an exponential family of distributions $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ with mean $\mbox{E}(y_{i}|\boldsymbol{u})=\mu_{i}$ and dispersion parameter $\phi>0$; the latter appears in the conditional variance of the distribution, $\mbox{Var}(y_{i}|\boldsymbol{u})=\phi V(\mu_{i})$.
@@
Linear mixed models arise as a special case of (1) when $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ is a normal distribution, $g(\boldsymbol{\mu})=\boldsymbol{\mu}$ is set to the identity link, and $\phi$ is the error variance such that $V(\mu)=1$. Two prominent non-normal cases of GLMMs are binomial GLMMs where $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ is the binomial distribution, $g(\mu)=\log\{\mu/(1-\mu)\}$ is the canonical logit link, $\phi=1$ is known and $V(\mu)=\mu(1-\mu)$, and Poisson GLMMs where $p(y_{i}|\boldsymbol{u};\boldsymbol{\beta},\phi)$ is the Poisson distribution, $g(\mu)=\log(\mu)$ is the canonical log link, $\phi=1$ is known and $V(\mu)=\mu$. The above formulation can be further expanded to involve more than one dispersion parameter, e.g., for Tweedie responses, and the REML methods reviewed in this article can be modified to accommodate such settings. However, for ease of presentation we will focus on the above formulation and refer the reader to McCullagh and Nelder (1989) and Wood (2017) for general overviews on exponential families of distributions for regression modeling.
@@
We assume the random effects in (1) are drawn from a multivariate normal distribution with a mean vector of zeros, and a random effects covariance matrix $\mbox{Cov}(\boldsymbol{u})=\boldsymbol{D}(\boldsymbol{\theta})$ which depends on a set of parameters $\boldsymbol{\theta}$. While it is possible to consider non-normally distributed random effects (see Section 7 for a discussion on REML estimation in HGLMs), we restrict the focus of this review largely to the normal random effects case given this is the most prevalent assumption in practice (Breslow and Clayton, 1993; McCulloch and Searle, 2004; Stroup, 2012). Note $\boldsymbol{D}(\boldsymbol{\theta})$ can and often does possess a large amount of structure, e.g., it may have a block-diagonal or fully diagonal structure in the case of independent cluster GLMMs. The diagonal elements of $\boldsymbol{D}(\boldsymbol{\theta})$ are often referred to as variance components, although in this article we will use the terms “variance components” and “random effects covariance matrix” interchangeably to mean $\boldsymbol{\theta}$ or $\boldsymbol{D}(\boldsymbol{\theta})$.
@@
Model (1) is broad enough to accommodate many applications of GLMMs. For instance, for independent clustered, nested or multi-level GLMMs, the columns of $\boldsymbol{Z}$ can be rearranged so that $\boldsymbol{Z}=[\boldsymbol{Z}_{1}\ldots\boldsymbol{Z}_{K}]$, with all the submatrices $\boldsymbol{Z}_{k}$, $k=1,\ldots,K$, being block-diagonal in form, while $\boldsymbol{D}(\boldsymbol{\theta})$ is also written in a block-diagonal form. Crossed-effect GLMMs are a case where the submatrices $\boldsymbol{Z}_{k}$ are not simultaneously block-diagonal. In this article, we will as much as possible utilize this broad formulation of GLMMs to keep the review general, with specific cases used as necessary, e.g., when particular papers develop REML estimation for a specific case of (1), and also in our simulation study.
@@
Standard unrestricted maximum likelihood estimation is based on optimizing equation (2) or some variation of this such as approximated marginal log-likelihoods obtained via Laplace, quadrature, variational, and Monte-Carlo approximations; see for example, Breslow and Clayton (1993), Shun and McCullagh (1995), McCulloch (1997), Ormerod and Wand (2012), Hall et al. (2020), and Hui (2021) and references therein. Often central to fitting GLMMs is the joint log-likelihood function of the responses and random effects
$\begin{split}\ell_{J}(\boldsymbol{\beta},\phi,\boldsymbol{\theta};\boldsymbol{y},\boldsymbol{u})=\log\{p(\boldsymbol{y}|\boldsymbol{u};\boldsymbol{\beta},\phi)\}\\ +\log\{p(\boldsymbol{u};\boldsymbol{\theta})\};\end{split}$ (3)
in particular the Laplace approximation is implemented around the maximizer of the joint likelihood function (Breslow and Clayton, 1993; Shun and McCullagh, 1995).